OpenStudy (anonymous):

Acetic acid contains carbon,hydogen and oxygen. 4324g of sample acetic acid on complete combustion gave 6.21g of CO_2 and 2.54g of H_2O. The Molecular mass of acetic acid is 60. Find (i) Empirical formula of acetic acid (ii) Molecular formula of acetic acid

sam (.sam.):

Can you draw a table for finding empirical formula?

OpenStudy (anonymous):

Step 1 = % of elements Step 2 =Mole Ratio step 3 = Simplest Ratio Step 4 = Establishing emp. formula

OpenStudy (anonymous):

Although i am good at this........but this question is crazy....cant just get it done

OpenStudy (anonymous):

Example Question : 1.8g of an organic compound on combustion gave 3.21g CO2 AND 1.322g H2O. Calculate Empirical Formula. Step 1 % of elements : C= 12/44 * 3.21/1.8 * 100 = 48.63% H=2/18*1.322/1.8 * 100 = 8.16 % O = 100-(48.63+8.16)=43.21% Step 2 Mole Ratio : C= 48.63/12 = 4.05 H= 8.16/1 = 8.16 O= 23.21/16 = 2.7 Step 3 Simplest Ratio : C=4.05/2.7 = 1.5 H=8.16/2.7 = 3 O=2.7/2.7 In order to get nearest whole no. ratio,all are multiplied by 2 Hence C=3 , H=6 , O=2 Step 4 Establishing Empirical Formula: C_3H_6_O_2

OpenStudy (anonymous):

^here is a example

OpenStudy (anonymous):

but the question i've asked is similar to this but it is entirely different ...it is giving crazy simplest ratio

sam (.sam.):

Hold on

OpenStudy (anonymous):

okay

OpenStudy (mos1635):

4324g of sample acetic...... is the number correct?????

OpenStudy (anonymous):

yes it is

OpenStudy (mos1635):

1,69 gr C 0.28 gr H 4324-1,69-0.28=4322.03 gr of oxyxen !!!!! something is wrong!!!

OpenStudy (anonymous):

it is from a question paper....and it is correct

OpenStudy (mos1635):

is CO2 and H2O in gr or in Kgr ????

OpenStudy (anonymous):

it is the 1st question i've asked

OpenStudy (mos1635):

i see it. i still thing the numbers are wrong. try to take CO2 and H2O in Krg to see if it make sense....

OpenStudy (anonymous):

4.324 kg acetic acid will produce only 6.21 g of CO2 and 2.54 H2O . Not possible, I think.

OpenStudy (anonymous):

okay....this is what i think since you know the compound is acetic acid, then you should know it has a molecular formula C2H4O2 and an empirical formula of CH2O.....